# Competent Discrete Time Modeling For analogue controlled PWM Converter Considering State-Feedback

### v2

@@ V-C Fixed-Frequency Trailing-Edge (Peak) PWM
In trailing-edge fixed-frequency PWM, the clock forces turn-on at the beginning of each switching period, while the comparator triggers turn-off (peak) within the period. The fixed-frequency constraint is
$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s}.$ (20)
To use the event equation at the sampling instant, sampling is chosen at the peak (turn-off) edge. A peak-to-peak map spans across the cycle boundary and depends on $T_{\mathrm{on},k}$ and $T_{\mathrm{on},k+1}$.
A. Large-signal peak-to-peak map
Let $x_{k}$ be sampled at the peak of cycle $k$. Then
$\text{off in cycle }k:\ T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k},\quad\text{then on in cycle }k{+}1:\ T_{\mathrm{on},k+1}.$
Hence
$x_{k+1}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on},k+1})\Big(\Phi_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})x_{k}$ (21)
$\qquad\qquad+\Gamma_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})\Big)+\Gamma_{\mathrm{on}}(T_{\mathrm{on},k+1}).$
B. Jacobian blocks with two timing perturbations
At steady state $T_{\mathrm{on}}^{\star}$ and $T_{\mathrm{off}}^{\star}=T_{s}-T_{\mathrm{on}}^{\star}$, define
$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (22)
$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
@@ B. Jacobian blocks with two timing perturbations
Define two timing sensitivities:
@@
$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}^{\star}x^{\star}+\Gamma_{\mathrm{off}}^{\star}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U,$
and, by the chain rule $\partial(T_{s}-T_{\mathrm{on},k})/\partial T_{\mathrm{on},k}=-1$,
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}$ (25)
@@
C. Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
At the peak sampling instant, $\Delta T=T_{\mathrm{on},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
@@ C. Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{on},k}$ (28)
$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
The appearance of $\widehat{T}_{\mathrm{on},k}$ is intrinsic to fixed-frequency peak sampling and maps to a $z^{-1}$ factor in the $z$-domain formulation.
@@ V-D Fixed-Frequency Leading-Edge (Valley) PWM
In leading-edge fixed-frequency PWM, the clock forces turn-off at the beginning of each period, while the comparator triggers turn-on (valley) within the period. The fixed-frequency constraint (20) holds. Sampling is taken at the valley (turn-on) edge, yielding a valley-to-valley map that depends on $T_{\mathrm{off},k}$ and $T_{\mathrm{off},k+1}$.
A. Large-signal valley-to-valley map
Let $x_{k}$ be sampled at the valley of cycle $k$. Then
$\text{on in cycle }k:\ T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k},$
$\text{then off in cycle }k{+}1:\ T_{\mathrm{off},k+1}.$
Hence
$x_{k+1}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off},k+1})\Big(\Phi_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})x_{k}$ (29)
$\qquad\qquad+\Gamma_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})\Big)+\Gamma_{\mathrm{off}}(T_{\mathrm{off},k+1}).$
B. Jacobian blocks with two timing perturbations
At steady state $T_{\mathrm{off}}^{\star}$ and $T_{\mathrm{on}}^{\star}=T_{s}-T_{\mathrm{off}}^{\star}$, define
$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (30)
$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
$\Gamma_{\mathrm{on}}^{\star}$ $=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$
@@ B. Jacobian blocks with two timing perturbations
Define
@@
$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}^{\star}x^{\star}+\Gamma_{\mathrm{on}}^{\star}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U,$
and, using $\partial(T_{s}-T_{\mathrm{off},k})/\partial T_{\mathrm{off},k}=-1$,
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}$ (33)
@@
C. Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
At the valley sampling instant, $\Delta T=T_{\mathrm{off},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
@@ C. Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{off},k}$ (36)
$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
As in the trailing-edge case, the previous-cycle timing term $\widehat{T}_{\mathrm{off},k}$ maps to a $z^{-1}$ factor in the $z$-domain representation.
@@ C. Fixed-frequency PWM: only one edge moves (no future accumulation)
VII Li–Jian Distillation: Port-Structured Weak-Coupling Approximation and MIMO Balances
This section presents the Li–Jian distillation procedure, i.e., a matrix-based reduction that converts a fully coupled PWL state-space model into a port-structured MIMO form whose kernel is an integrator cascade. The reduced model (i) preserves the physically chosen output/port variable (through a known $C$-row), (ii) replaces internal resistive couplings by controlled-source cascades driven by ports, and (iii) exposes volt-second and amp-second balance as solvability / periodicity conditions of the distilled Poincaré map.

### v3

@@ V-C Fixed-Frequency Trailing-Edge (Peak) PWM
In fixed-frequency trailing-edge PWM, the switch is turned on by the clock at the beginning of each cycle, while the comparator triggers turn-off within the cycle. The switching period is constant, hence for every cycle $k$,
$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s},\qquad T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k}.$ (20)
Let the ramp have amplitude $V_{r}$ over one period $T_{s}$, so its slope is
$S_{e}:=\frac{V_{r}}{T_{s}}.$ (21)
V-C1 Large-signal peak-to-peak map
Let $x_{k}$ be sampled at the peak (turn-off) edge of cycle $k$. From this sampling instant to the next peak, the trajectory consists of (i) the off-interval of cycle $k$ with duration $T_{\mathrm{off},k}$, followed by (ii) the on-interval of cycle k+1k{+}1 with duration $T_{\mathrm{on},k+1}$. Therefore,
$x_{k+1}$ $=\Phi_{\mathrm{on}}\!\big(T_{\mathrm{on},k+1}\big)\Big(\Phi_{\mathrm{off}}\!\big(T_{\mathrm{off},k}\big)x_{k}+\Gamma_{\mathrm{off}}\!\big(T_{\mathrm{off},k}\big)\Big)$ (22)
$\quad+\Gamma_{\mathrm{on}}\!\big(T_{\mathrm{on},k+1}\big),$
where, for each mode $m\in\{\mathrm{on},\mathrm{off}\}$, $\Phi_{m}(T)=e^{A_{m}T}$ and $\Gamma_{m}(T)=\int_{0}^{T}e^{A_{m}(T-\tau)}B_{m}U\,d\tau$.
V-C2 Jacobian blocks with two timing perturbations
At the steady state, denote
$T_{\mathrm{on}}^{\star},\quad T_{\mathrm{off}}^{\star}:=T_{s}-T_{\mathrm{on}}^{\star},\qquad x^{\star}.$ (23)
Define
$\Phi_{\mathrm{on}}^{\star}:=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),\quad\Phi_{\mathrm{off}}^{\star}:=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),\quad\Gamma_{\mathrm{off}}^{\star}:=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}^{\star}).$ (24)
@@ V-C2 Jacobian blocks with two timing perturbations
Define two timing sensitivities corresponding to the two durations that appear in (22). First, the sensitivity to the next-cycle on-time $T_{\mathrm{on},k+1}$:
@@
$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}^{\star}x^{\star}+\Gamma_{\mathrm{off}}^{\star}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U.$
Second, the sensitivity to the current-cycle on-time $T_{\mathrm{on},k}$, which enters (22) only through $T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k}$ in (20):
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}=-\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}$ (27)
@@
V-C3 Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
At the peak (turn-off) sampling instant, the edge time is $\Delta T=T_{\mathrm{on},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$. Using the standard ramp-comparator small-signal relation,
@@ V-C3 Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{on},k}-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$ (30)
@@ V-D Fixed-Frequency Leading-Edge (Valley) PWM
In fixed-frequency leading-edge PWM, the clock forces turn-off at the beginning of each cycle, while the comparator triggers turn-on (valley) within the cycle. The fixed-period constraint implies, for every cycle $k$,
$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s},\qquad T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}.$ (31)
The ramp slope $S_{e}$ is still given by (21).
V-D1 Large-signal valley-to-valley map
Let $x_{k}$ be sampled at the valley (turn-on) edge of cycle $k$. From this sampling instant to the next valley, the trajectory consists of (i) the on-interval of cycle $k$ with duration $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$, followed by (ii) the off-interval of cycle k+1k{+}1 with duration $T_{\mathrm{off},k+1}$. Therefore,
$x_{k+1}$ $=\Phi_{\mathrm{off}}\!\big(T_{\mathrm{off},k+1}\big)\Big(\Phi_{\mathrm{on}}\!\big(T_{\mathrm{on},k}\big)x_{k}+\Gamma_{\mathrm{on}}\!\big(T_{\mathrm{on},k}\big)\Big)$ (32)
$\quad+\Gamma_{\mathrm{off}}\!\big(T_{\mathrm{off},k+1}\big),$
with $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$ per (31).
V-D2 Jacobian blocks with two timing perturbations
At the steady state, denote
$T_{\mathrm{off}}^{\star},\quad T_{\mathrm{on}}^{\star}:=T_{s}-T_{\mathrm{off}}^{\star},\qquad x^{\star}.$ (33)
Define
$\Phi_{\mathrm{on}}^{\star}:=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),\quad\Phi_{\mathrm{off}}^{\star}:=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),\quad\Gamma_{\mathrm{on}}^{\star}:=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$ (34)
@@ V-D2 Jacobian blocks with two timing perturbations
Define the two timing sensitivities. First, the sensitivity to the next-cycle off-time $T_{\mathrm{off},k+1}$:
@@
$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}^{\star}x^{\star}+\Gamma_{\mathrm{on}}^{\star}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U.$
Second, the sensitivity to the current-cycle off-time $T_{\mathrm{off},k}$ which enters (32) only through $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$ in (31):
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}=-\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}$ (37)
@@
V-D3 Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
At the valley (turn-on) sampling instant, the edge time is $\Delta T=T_{\mathrm{off},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$. Thus
@@ V-D3 Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{off},k}-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$ (40)
@@ C. Fixed-frequency PWM: only one edge moves (no future accumulation)
VII Jian–Li Distillation: Port-Structured Weak-Coupling Approximation and MIMO Balances
This section presents the Jian–Li distillation procedure, i.e., a matrix-based reduction that converts a fully coupled PWL state-space model into a port-structured MIMO form whose kernel is an integrator cascade. The reduced model (i) preserves the physically chosen output/port variable (through a known $C$-row), (ii) replaces internal resistive couplings by controlled-source cascades driven by ports, and (iii) exposes volt-second and amp-second balance as solvability / periodicity conditions of the distilled Poincaré map.